# Pose2Gaze: Generating Realistic Human Gaze Behaviour from Full-body Poses using an Eye-body Coordination Model

### v1

@@ Abstract
While generating realistic body movements, e.g., for avatars in virtual reality, is widely studied in computer vision and graphics, the generation of eye movements that exhibit realistic coordination with the body remains under-explored. We first report a comprehensive analysis of the coordination of human eye and full-body movements during everyday activities based on data from the MoGaze and GIMO datasets. We show that eye gaze has strong correlations with head directions and also full-body motions and there exists a noticeable time delay between body and eye movements. Inspired by the analyses, we then present Pose2Gaze – a novel eye-body coordination model that first uses a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and then applies a convolutional neural network to generate realistic eye movements. We compare our method with state-of-the-art methods that predict eye gaze only from head movements for three different generation tasks and demonstrate that Pose2Gaze significantly outperforms these baselines on both datasets with an average improvement of $26.4\%$ and $21.6\%$ in mean angular error, respectively. Our findings underline the significant potential of cross-modal human gaze behaviour analysis and modelling.
[…]
1 Introduction
With the increasing use of augmented and virtual reality (VR)[36], the need to create virtual agents that behave realistically has grown significantly in recent years. Virtual agents that behave like humans is crucial for a variety of applications, such as online learning [31, 32], virtual interviewing and counselling [2, 7], virtual social interactions [40, 29], as well as large-scale virtual worlds [36]. Creating realistic behaviour is also important in character animation, e.g., in films [13]. However, previous work has mainly focused on generating realistic body movements [3, 28] despite the fact that eye movements are equally, if not even more, important in human-human encounters [41] and, as such, also for reducing the uncanny valley effect [42].
@@ 1 Introduction
Generating realistic eye movement behaviour is challenging given the highly variable spatio-temporal dynamics of gaze allocation as well as various top-down influences, e.g., of the visual world, the agent’s tasks or goals, or social norms that agents have ideally adhere to during interactions with humans [5]. Research on eye gaze prediction has revealed that eye movements are carefully coordinated with head movements and, as such, that head orientation can be used as a proxy to eye gaze direction [45, 21, 20, 18]. Consequently, previous works have mainly focused on exploiting eye-head coordination for generating eye movements [21, 20, 18].
In this work, we provide a complementary perspective and study, for the first time, eye-body coordination during everyday activities and explore the generation of realistic human gaze behaviour from full-body poses.
To this end we first report comprehensive analyses on the correlations between gaze direction and full-body movements on two publicly available datasets: MoGaze [26] and GIMO [51]. These datasets contain human eye and full-body movement data recorded using a body tracking system and a mobile eye tracker during various daily activities.To the best of our knowledge, we are first to study the coordination of eye movements and full-body movements simultaneously. Our analyses confirm that eye gaze direction is strongly correlated with head direction but also with the motions of different human body joints. We also find that there exists a noticeable time delay between full-body motions and eye movements.
@@
Based on these findings we present Pose2Gaze – the first method to generate realistic human gaze behaviour from human full-body poses. At the core of our method is a novel learning-based eye-body coordination model that first uses a convolutional neural network (CNN) and a spatio-temporal graph convolutional neural network (GCN) to extract features from head directions and full-body poses respectively and then employs a convolutional neural network to generate realistic human eye gaze from the extracted features. We compare our method with state-of-the-art gaze prediction methods for three different generation tasks, i.e. generating target eye gaze from past, present, and future body poses respectively, on the MoGaze and GIMO datasets and show that our method outperforms these baselines by a large margin on both the MoGaze ($26.4\%$ improvement) and GIMO datasets ($21.6\%$) in terms of mean angular gaze error [^1] Source code and trained models will be released upon acceptance..
The specific contributions of our work are three-fold:
  • We provide a comprehensive analysis of eye-body coordination in various daily activities and reveal that eye gaze is strongly correlated with head directions and also full-body motions.
@@
  • We propose Pose2Gaze, a novel eye-body coordination model that applies a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head directions and full-body poses respectively and uses a convolutional neural network to generate realistic eye gaze.
@@
  • We report extensive experiments on two public datasets for three different generation tasks and demonstrate significant performance improvements over several state-of-the-art methods.
2 Related Work
2.1 Human Movement Generation
@@
Generating realistic human movements for virtual agents is an important research topic in the area of virtual reality. Some researchers focused on generating human body movements from speech signals. For example, Hasegawa et al. used recurrent neural networks (RNNs) to generate natural human full-body poses from perceptual features extracted from the input speech audio[15] while Kucherenko et al. employed autoencoder neural networks to learn latent representations for human poses and speech signals and then learned the mappings between the two representations to generate human body motions[27]. Other researchers used the text transcripts of speech to generate human poses. Specifically, Yoon et al. proposed to use a RNN-based encoder to extract features from speech text and a RNN-based decoder to generate human poses[50] while Bhattacharya et al. employed an end-to-end transformer network to generate human movements from the text transcripts of speech[3]. In addition, there also exist some works that generated human dance motions from input music and achieved good performances[49, 30]. While previous works typically focused on generating human body movements, they neglected the generation of human eye gaze, which is significant for human-human[16, 8] and human-computer interactions[9, 44, 23]. In this work, we focus on generating realistic human eye gaze directly from human full-body poses to fill this gap.
2.2 Eye-body Coordination
@@
The coordination of human eye and body movements is an important research topic in the areas of cognitive science and human-centred computing. Many researchers focused on the coordinated movements between the eyes and the head[21, 46, 12]. Specifically, Stahl analysed the process of gaze shifts and found its amplitude to be proportional to that of head movements[46]. Fang et al. further examined eye-head coordination during visual search in a large visual field and found that eye-head coordination plays an important role in visual cognitive processing[12]. Hu et al. revealed that human eye movements are strongly correlated with head movements in both free-viewing[21, 20] and task-oriented settings[18, 19] in immersive virtual environments while Kothari et al. identified the coordination of eye and head movements in real-world daily activities[24]. Recently, some researchers went beyond eye-head coordination to investigate the correlations between eye movements and the movements of different body parts. For example, Sidenmark et al. focused on the gaze shift process in immersive virtual reality and discovered general eye, head, and torso coordination patterns[43]. Batmaz et al. developed a VR training system based on the coordination of eye and hand movements and compared user performance in three different virtual environments[1]. Emery et al. collected a large-scale dataset that contains human eye, hand, and head movements in a virtual environment and identified the coordination of eye, hand, and head motions[11]. Randhavane et al. investigated the effectiveness of eye-gait coordination on expressing emotions and further employed gaze and gait features to generate various emotions for virtual agents[39]. However, prior works typically focused on the correlations between eye gaze and a particular body part, e.g. head, hand, or torso. In contrast, we are the first to study the coordinations of eye and full-body movements simultaneously, which paves the way for generating realistic eye gaze from full-body poses.
2.3 Eye Gaze Prediction
@@
Eye gaze prediction is a popular research topic in the areas of computer vision and human-centred computing. Typical gaze prediction methods can be classified into bottom-up and top-down approaches. Bottom-up methods predict eye gaze from the low-level image features of the scene content such as intensity, colour, and orientation[22, 6]. For example, Itti et al. used multiscale colour, intensity, and orientation features extracted from the image to predict saliency map (density map of eye gaze distribution)[22] while Cheng et al. employed the global contrast features of the image to generate saliency map[6]. Top-down approaches take high-level features of the scene such as specific tasks and context information into consideration to predict eye gaze[4, 25]. For example, Borji et al. employed players’ input such as 2D mouse positions and joystick buttons to predict their eye gaze[4] while Koulieris et al. used game state variables to predict users’ eye gaze positions in video games[25]. In addition to the typical bottom-up and top-down approaches, recently some researchers took the eye-head coordination into consideration and attempted to predict eye gaze from human head movements[45, 21]. Specifically, Nakashima et al. proposed to use head directions as prior knowledge to improve the accuracy of bottom-up saliency prediction methods through simple multiplication of the predicted saliency map by a Gaussian head direction bias[38]. Sitzmann et al. employed users’ head orientations to predict saliency maps for 360-degree images and achieved an accuracy that is on par with the performance of bottom-up saliency predictors[45]. Hu et al. proposed to use users’ head rotation velocities to predict users’ eye gaze positions in immersive virtual environments and achieved good performances in both free-viewing[21, 20] and task-oriented situations[18, 17]. However, existing works have only explored the effectiveness of head movements on the task of eye gaze prediction. In stark contrast, in this work we focus on the coordination of eye and full-body movements and investigate the effectiveness of full-body movements on generating realistic eye gaze.
3 Analysis of Eye-body Coordination
3.1 Gaze and Motion Data
To study the coordination between human eye gaze and full-body movements, we conducted a comprehensive analysis based on the only public motion capture datasets that also offer eye gaze information: MoGaze [26] and GIMO [51]. The MoGaze dataset contains eye gaze and full-body poses of $21$ human joints from six people performing everyday pick and place activities in an indoor environment. GIMO dataset offers eye gaze and full-body poses of $23$ human joints from $11$ people performing daily activities in various indoor scenes. The activities include resting (sitting or laying on objects), interacting with objects (touching, holding, stepping on, reaching to objects), and changing the state of objects (opening, pushing, transferring, throwing, picking up, lifting, connecting, screwing, grabbing, swapping objects).
3.2 Correlation between Eye Gaze and Body Orientations
We used the forward directions of human body joints available in the datasets to represent body orientations. Specifically, we represented body orientations in MoGaze using the forward directions of head, neck, torso, pelvis, and base and indicated body orientations in GIMO using the directions of head, neck, spine3, spine2, spine1, and pelvis. We used the cosine similarity to analyse the correlation between eye gaze and body orientations. Cosine similarity measures the similarity between two non-zero vectors by calculating the cosine of the angle between the vectors and produces a value in the range from $-1$ to $+1$, where $-1$ indicates perfect negative correlation, 00 denotes no correlation, and $+1$ represents perfect positive correlation. Table 1 summarises the cosine similarities between eye gaze and body orientations in the MoGaze and GIMO datasets. As the table shows, in both datasets eye gaze direction is strongly correlated with the directions of different body joints. In particular, eye gaze exhibits very high correlation with head direction, achieving a cosine similarity of $0.92$ in MoGaze and $0.93$ in GIMO. We also find that the directions of different body joints have very strong correlations with each other, e.g. neck and head directions have a cosine similarity of $0.95$ in MoGaze and $0.92$ in GIMO, indicating that there exists redundant information among these directions. Therefore, for simplicity, we suggest to use only head direction to generate human eye gaze rather than using all the body directions.
Gaze Head Neck Torso Pelvis Base
Gaze 1 0.92 0.84 0.64 0.59 0.63
Head - 1 0.95 0.81 0.78 0.81
Neck - - 1 0.94 0.91 0.93
Torso - - - 1 0.99 0.99
Pelvis - - - - 1 0.99
Base - - - - - 1
(a) Gaze-body orientations in MoGaze
Gaze Head Neck Spine3 Spine2 Spine1 Pelvis
Gaze 1 0.93 0.89 0.84 0.77 0.84 0.72
Head - 1 0.92 0.84 0.77 0.84 0.71
Neck - - 1 0.97 0.95 0.97 0.90
Spine3 - - - 1 0.98 0.99 0.94
Spine2 - - - - 1 0.97 0.97
Spine1 - - - - - 1 0.93
Pelvis - - - - - - 1
(b) Gaze-body orientations in GIMO
@@ 3.2 Correlation between Eye Gaze and Body Orientations
Table 1: The cosine similarities between eye gaze directions and the directions of different body joints in the MoGaze and GIMO datasets. Gaze directions are strongly correlated with body orientations, especially with head directions. The directions of different body joints have strong correlations with each other.
@@
To investigate whether there exist time delays between head direction and eye gaze, we added different time intervals between head and gaze directions and further calculated their cosine similarities. We can see from Figure 1 that head direction achieves its highest correlation with eye gaze at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, indicating that there exists little or no time delay between head and eye movements.
(a) Gaze-head direction in MoGaze
(b) Gaze-head direction in GIMO
@@
Figure 1: The cosine similarities between head and gaze directions at different time intervals in the (a) MoGaze and (b) GIMO datasets. Time interval refers to the time difference between head and gaze directions. The highest gaze-head correlation occurs at $100$ $ms$ in MoGaze and 00 $ms$ in GIMO, suggesting that there is little or no time delay between head and eye movements.
3.3 Correlations between Eye Gaze and Body Movements
To analyse the correlations between human eye gaze and human full-body movements, we calculated the velocities of different body joints and normalised these velocities to 3D unit vectors to represent the directions of body motions. We further calculated the cosine similarities between eye gaze and the directions of body motions to measure their correlations. Figure 2 shows the cosine similarities between eye gaze and full-body motions in the MoGaze and GIMO datasets, respectively. We can see that eye gaze directions have strong correlations with the motions of different body joints in both datasets.
(a) Gaze-body motions in MoGaze
(b) Gaze-body motions in GIMO
Figure 2: The cosine similarities between eye gaze directions and the directions of body motions in the (a) MoGaze and (b) GIMO datasets. Eye gaze is strongly correlated with human full-body motions.
To investigate any potential time delays between body motions and eye movements, we added different time intervals between body motions and gaze directions and further calculated their correlations. Specifically, for simplicity we first grouped all the human joints into three different body parts, i.e. Torso, Arm, and Leg. In the MoGaze dataset Torso consists of the body joints of base, pelvis, torso, neck, and head while in the GIMO dataset Torso contains the body joints of pelvis, spine1, spine2, spine3, neck, head, and jaw. In both datasets, Arm contains the body joints of collar, shoulder, elbow, and wrist while Leg includes hip, knee, ankle, and foot, where the left and right joints are averaged into one single joint, e.g. left shoulder and right shoulder are represented using shoulder. We then calculated the cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific joints. The correlation between gaze and a body part was computed using the mean of the cosine similarities between eye gaze and the specific joints within the body part. We can see from Figure 3 that in both grouped body parts and specific body joints, the highest correlations consistently occur between current eye gaze and body motions in the near future, suggesting that there exists a noticeable time delay between body motions and eye movements.
(a) Gaze-body motions in MoGaze
(b) Gaze-torso motions in MoGaze
(c) Gaze-arm motions in MoGaze
(d) Gaze-leg motions in MoGaze
(e) Gaze-body motions in GIMO
(f) Gaze-torso motions in GIMO
(g) Gaze-arm motions in GIMO
(h) Gaze-leg motions in GIMO

### v3

# Pose2Gaze: Eye-body Coordination during Daily Activities for Gaze Prediction from Full-body Poses

@@ Abstract
Human eye gaze plays a significant role in many virtual and augmented reality (VR/AR) applications, such as gaze-contingent rendering, gaze-based interaction, or eye-based activity recognition. However, prior works on gaze analysis and prediction have only explored eye-head coordination and were limited to human-object interactions. We first report a comprehensive analysis of eye-body coordination in various human-object and human-human interaction activities based on four public datasets collected in real-world (MoGaze), VR (ADT), as well as AR (GIMO and EgoBody) environments. We show that in human-object interactions, e.g. pick and place, eye gaze exhibits strong correlations with full-body motion while in human-human interactions, e.g. chat and teach, a person’s gaze direction is correlated with the body orientation towards the interaction partner. Informed by these analyses we then present Pose2Gaze – a novel eye-body coordination model that uses a convolutional neural network and a spatio-temporal graph convolutional neural network to extract features from head direction and full-body poses, respectively, and then uses a convolutional neural network to predict eye gaze. We compare our method with state-of-the-art methods that predict eye gaze only from head movements and show that Pose2Gaze outperforms these baselines with an average improvement of $24.0\%$ on MoGaze, $10.1\%$ on ADT, $21.3\%$ on GIMO, and $28.6\%$ on EgoBody in mean angular error, respectively. We also show that our method significantly outperforms prior methods in the sample downstream task of eye-based activity recognition. These results underline the significant information content available in eye-body coordination during daily activities and open up a new direction for gaze prediction.
Index Terms:
Eye-body coordination, human-object interaction, human-human interaction, gaze prediction, activity recognition, virtual reality, augmented reality
I Introduction
With the increasing use of virtual and augmented reality (VR/AR), understanding and predicting human behaviour in VR/AR environments has become a popular research topic and is a key requirement for intelligent human-aware VR/AR agents [1, 2]. Human eye movements are of particular interest given that they are key for important VR/AR applications, such as 1) gaze-contingent rendering that renders the gaze central region with high quality while decreasing the fidelity in peripheral region to increase rendering efficiency [3]; 2) gaze-based interaction that uses eye gaze to interact with 3D objects in virtual environments [4]; 3) virtual content optimisation that optimises the design of virtual content based on gaze distributions [5]; 4) gaze-guided redirected walking that redirects a user’s walking path based on their gaze directions [6]; or 5) eye-based activity recognition that recognises user activities from their eye movements [7].
@@ I Introduction
Understanding and predicting human eye movements in VR/AR is challenging given the highly variable spatio-temporal dynamics of gaze allocation as well as various top-down influences, e.g. specific tasks that a user has to perform [1, 7] or social norms that a user ideally adheres to during interactions with humans [8]. Research on gaze behaviour analysis has revealed that eye movements are closely coordinated with head movements and, as such, that head orientation can be used as a proxy to eye gaze [9, 10, 11, 1]. Consequently, previous works have mainly focused on exploiting eye-head coordination for predicting eye movements [10, 11, 1] while neglecting coordination between eye and full-body movements. Furthermore, most existing works have studied human-object interactions [1, 7] and have neglected the more challenging but for VR/AR applications also more practically useful human-human interactions, in which a person’s eye gaze is influenced by the body movements of the interaction partner.
@@
To fill this gap, in this work we are the first to report comprehensive analyses of the correlations between gaze direction and full-body movements using four publicly available datasets collected in the real world (MoGaze [12]), in VR (ADT [13]) and AR (GIMO [14] and EgoBody [15]). These datasets contain human eye and full-body motion data recorded using a body tracking system and an eye tracker during various daily activities. Our analyses reveal that in human-object interactions, e.g. pick, place, touch, or hold, eye gaze exhibits strong correlations with full-body motion and eye movements precede body movements. In contrast, in human-human interactions, e.g. chat, teach, or discuss, a person’s gaze is more closely correlated with the body orientation towards the interaction partner. Based on these findings we further present Pose2Gaze – the first method to predict eye movements from human full-body poses during human-object and human-human interactions. At the core of our method is a novel learning-based eye-body coordination model that based on a convolutional neural network (CNN) and a spatio-temporal graph convolutional neural network (GCN) to extract features from head directions and full-body poses, respectively. A convolutional neural network is also used to predict human eye gaze from these extracted features. We compare our method with state-of-the-art gaze prediction methods for three different prediction tasks, i.e. generating target eye gaze from past, present, and future body poses, respectively, and show that our method outperforms these baselines by a large margin in terms of mean angular error on the MoGaze ($24.0\%$ improvement), ADT ($10.1\%$), GIMO ($21.3\%$), and EgoBody ($28.6\%$) datasets. We also evaluate the effectiveness of our method for a sample downstream task, i.e eye-based activity recognition, and demonstrate that using our gaze prediction method results in significant improvements in recognition accuracy. The full source code and trained models are available at zhiminghu.net/hu24_pose2gaze.
The specific contributions of our work are four-fold:
  • We provide comprehensive analyses of eye-body coordination in human-object and human-human interactions in the real world, VR, and AR, and show that eye gaze is strongly correlated with full-body motion in human-object interactions and is closely linked with the directions between two bodies in human-human interactions.
@@
  • We propose Pose2Gaze – a novel eye-body coordination model for gaze prediction that combines a CNN and a spatio-temporal GNN to extract features from head directions and full-body poses.
@@
  • We report extensive experiments on four public datasets for three different prediction tasks and demonstrate significant performance improvements over several state-of-the-art methods.
  • We demonstrate our method’s effectiveness in the sample downstream task of eye-based activity recognition.
II Related Work
II-A Human Movement Prediction
@@
Human movement prediction is an important research topic in the areas of virtual reality and augmented reality. Some researchers focused on predicting human body movements from speech signals. For example, Hasegawa et al. used recurrent neural networks (RNNs) to generate natural human full-body poses from perceptual features extracted from the input speech audio[16] while Kucherenko et al. employed autoencoder neural networks to learn latent representations for human poses and speech signals and then learned the mappings between the two representations to predict human body motions[17]. Other researchers used the text transcripts of speech to generate human poses. Specifically, Yoon et al. proposed to use a RNN-based encoder to extract features from speech text and a RNN-based decoder to generate human poses[18] while Bhattacharya et al. employed an end-to-end Transformer network to predict human movements from the text transcripts of speech[19]. In addition, there also exist some works that generated human dance motions from input music and achieved good performances[20, 21]. While previous works typically focused on generating human body movements, they neglected the prediction of human eye gaze, which is significant for human-human[22, 23] and human-computer interactions[24, 4, 25]. In this work, we focus on predicting human eye movements directly from human full-body poses to fill this gap.
II-B Eye-body Coordination
@@
The coordination of human eye and body movements is an important research topic in the areas of cognitive science and human-centred computing. Many researchers focused on the coordinated movements between the eyes and the head[10, 26, 27]. Specifically, Stahl analysed the process of gaze shifts and found its amplitude to be proportional to that of head movements[26]. Fang et al. further examined eye-head coordination during visual search in a large visual field and found that eye-head coordination plays an important role in visual cognitive processing[27]. Hu et al. revealed that human eye movements are strongly correlated with head movements in both free-viewing[10, 11] and task-oriented settings[1, 7] in immersive virtual environments while Kothari et al. identified the coordination of eye and head movements in real-world daily activities[28]. Recently, some researchers went beyond eye-head coordination to investigate the correlations between eye movements and the movements of different body parts. For example, Sidenmark et al. focused on the gaze shift process in immersive virtual reality and discovered general eye, head, and torso coordination patterns[29]. Batmaz et al. developed a VR training system based on the coordination of eye and hand movements and compared user performance in three different virtual environments [30]. Emery et al. collected a large-scale dataset that contains human eye, hand, and head movements in a virtual environment and identified the coordination of eye, hand, and head motions[31]. Randhavane et al. investigated the effectiveness of eye-gait coordination on expressing emotions and further employed gaze and gait features to generate various emotions for virtual agents[32]. However, prior works typically focused on the correlations between eye gaze and a particular body part, e.g. head, hand, or torso. In contrast, we are the first to study the coordinations of eye and full-body movements simultaneously, which paves the way for predicting eye gaze from full-body poses.
II-C Eye Gaze Prediction
@@
Eye gaze prediction is a popular research topic in the areas of computer vision and human-centred computing. Typical gaze prediction methods can be classified into bottom-up and top-down approaches. Bottom-up methods predict eye gaze from the low-level image features of the scene content such as intensity, colour, and orientation[33, 34]. For example, Itti et al. used multiscale colour, intensity, and orientation features extracted from the image to predict saliency map (density map of eye gaze distribution)[33] while Cheng et al. employed the global contrast features of the image to generate saliency map[34]. Top-down approaches take high-level features of the scene such as specific tasks and context information into consideration to predict eye gaze[35, 36]. For example, Borji et al. employed players’ input such as 2D mouse positions and joystick buttons to predict their eye gaze[35] while Koulieris et al. used game state variables to predict users’ eye gaze positions in video games[36]. In addition to the typical bottom-up and top-down approaches, recently some researchers took the eye-head coordination into consideration and attempted to predict eye gaze from human head movements[9, 10]. Specifically, Nakashima et al. proposed to use head directions as prior knowledge to improve the accuracy of bottom-up saliency prediction methods through simple multiplication of the predicted saliency map by a Gaussian head direction bias[37]. Sitzmann et al. employed users’ head orientations to predict saliency maps for 360-degree images and achieved an accuracy that is on par with the performance of bottom-up saliency predictors[9]. Hu et al. proposed to use users’ head rotation velocities to predict users’ eye gaze positions in immersive virtual environments and achieved good performances [1, 38]. However, existing works have only explored the effectiveness of head movements on the task of eye gaze prediction and were limited to human-object interactions. In stark contrast, in this work we investigate the effectiveness of full-body movements on predicting human eye gaze in both human-object and human-human interaction activities.
III Analysis of Eye-body Coordination
III-A Gaze and Pose Data
We studied the eye-body coordination from two perspectives, i.e. the correlation between gaze direction and body orientations and the correlation between gaze direction and body motions (translational movements of the body). We conducted a comprehensive analysis based on four public datasets that contain various human-object and human-human interaction activities collected from the real world (MoGaze [12]), VR environments (ADT [13]), as well as AR scenarios (GIMO [14] and EgoBody [15]).
MoGaze
The MoGaze dataset contains 180 minutes of full-body motion capture data collected from seven participants performing everyday pick and place activities in an indoor environment. One participant’s eye gaze data is not recorded. Therefore, we only used the data from six people to analyse eye-body coordination.
ADT
The ADT dataset collects human eye gaze and/or full-body pose data performing various indoor activities in two virtual environments (an apartment and an office environment). To analyse eye-body coordination, we used the $34$ sequences that contain both eye gaze and full-body pose data. Each sequence lasts for around 2 minutes and the activities include decoration, meal, and work.
GIMO
The GIMO dataset records eye gaze and full-body poses from $11$ people performing various daily activities in various indoor scenes. The whole dataset contains $215$ sequences and each sequence lasts for around 10 seconds. The activities can be classified into three categories, i.e. change the state of objects (open, push, transfer, throw, pick up, lift, connect, screw, grab, swap objects), interact with objects (touch, hold, step on, reach to objects), and rest (sit or lay on objects). The original dataset does not provide activity labels for each sequence. To gain a comprehensive analysis, we manually labelled each sequence as one of the three categories.
EgoBody
The EgoBody dataset collects eye gaze and full-body pose data from $36$ people performing diverse human-human interaction activities (between two humans) in 15 indoor scenes. The whole dataset contains $125$ sequences and each sequence lasts for around 2 minutes. The activities include catch, chat, dance, discuss, learn, perform, and teach.
III-B Correlation between Eye Gaze and Body Orientations
@@
TABLE I: The cosine similarities between eye gaze direction and the directions of different body joints in the MoGaze, GIMO, and EgoBody datasets. Gaze direction is strongly correlated with body orientations, especially with head direction.
base pelvis torso neck head
MoGaze pick 0.64 0.60 0.66 0.84 0.92
place 0.62 0.58 0.63 0.84 0.92
GIMO change 0.76 0.86 0.86 0.90 0.93
interact 0.72 0.82 0.83 0.87 0.93
rest 0.67 0.82 0.83 0.87 0.92
EgoBody catch 0.90 0.94 0.94 0.96 0.97
chat 0.81 0.85 0.87 0.90 0.94
dance 0.82 0.86 0.87 0.93 0.97
discuss 0.88 0.88 0.91 0.93 0.94
learn 0.70 0.75 0.77 0.84 0.89
perform 0.90 0.92 0.92 0.95 0.97
teach 0.84 0.84 0.86 0.89 0.93
Prior works have mainly focused on the correlation between eye gaze and head directions [10, 11]. To gain a comprehensive understanding of eye-body coordination, we simultaneously analysed the correlation between eye gaze and the orientations of different body joints, e.g. neck and torso. The ADT dataset does not provide the orientations of the body joints and therefore we only performed analysis on the MoGaze, GIMO, and EgoBody datasets. Specifically, we employed the forward directions of the body joints to represent body orientations and used the cosine similarity to analyse the correlation between eye gaze and body orientations. Cosine similarity measures the similarity between two non-zero vectors by calculating the cosine of the angle between the vectors and produces a value in the range from $-1$ to $+1$, where $-1$ indicates perfect negative correlation, 00 denotes no correlation, and $+1$ represents perfect positive correlation. Table I summarises the cosine similarities between eye gaze direction and the directions of different body joints in the MoGaze, GIMO, and EgoBody datasets. We can see that most of the cosine similarities are larger than $0.6$, demonstrating that eye gaze has very high correlations with body orientations in various daily activities. In particular, we find that in all the activities eye gaze consistently exhibits highest correlation with head direction ($>0.9$ in most cases), validating that head direction correlates better with eye gaze than other body orientations.
@@
To investigate whether there exist time delays between eye gaze and head orientations, we desynchronised the gaze and head signals by adding different time intervals between them (time interval $>0$ means head data is moved forward in time while time interval $<0$ means gaze data is moved forward) and further calculated their cosine similarities. We can see from Figure 1 that head direction achieves its highest correlation with eye gaze at the time interval of between $-100$ and $200$ $ms$, indicating that there exists little or no time delay between head and eye movements. These results further validate that head direction is a suitable proxy for eye gaze in everyday activities.
(a) MoGaze
(b) ADT
(c) GIMO
(d) EgoBody
@@
Figure 1: The cosine similarities between head and gaze directions at different time intervals in the (a) MoGaze (b) ADT (c) GIMO and (d) EgoBody datasets. The highest correlations occur at between $-100$ and $200$ $ms$, suggesting that there is little or no time delay between head and eye movements.
III-C Correlations between Eye Gaze and Body Motions
To analyse the correlations between human eye gaze and human full-body motion, we calculated the velocities of different body joints using the difference in position between two consecutive frames and then normalised these velocities to 3D unit vectors to represent the directions of body motions. We further calculated the cosine similarities between eye gaze and the directions of body motions to measure their correlations. Table II shows the cosine similarities between eye gaze and full-body motions in the MoGaze, ADT, GIMO, and EgoBody datasets. We find that in the human-object interaction activities, e.g. pick, place, decoration, change, and interact, eye gaze exhibits strong correlations with full-body motions ($>$ 0.3 in most cases). In the meal and work activities, the eye-body correlations are relatively smaller, probably because the body motions are less frequent in such activities and are thus less correlated with eye gaze. In the human-human interaction activities, e.g. chat, discuss, and learn, we find that eye gaze has little or no correlation with body motions ($\leq$ 0.05 in most cases). This is probably because a person’s eye and body motion are strongly influenced by the interaction partner in human-human interactions, thus degrading the eye-body correlations.
TABLE II: The cosine similarities between eye gaze and the motions of different body joints in the MoGaze, ADT, GIMO, and EgoBody datasets. Eye gaze has strong correlations with body motions in human-object interaction activities while having little or no correlation in human-human interaction activities. l_col: left collar, r_col: right collar, l_sho: left shoulder, r_sho: right shoulder, l_elb: left elbow, r_elb: right elbow, l_wri: left wrist, r_wri: right wrist, l_hip: left hip, r_hip: right hip, l_kne: left knee, r_kne: right knee, l_ank: left ankle, r_ank: right ankle, l_toe: left toe, r_toe: right toe.
base pelvis torso neck head l_col r_col l_sho r_sho l_elb r_elb l_wri r_wri l_hip r_hip l_kne r_kne l_ank r_ank l_toe r_toe Average
MoGaze pick 0.40 0.40 0.41 0.42 0.46 0.42 0.42 0.38 0.41 0.35 0.40 0.34 0.46 0.40 0.40 0.42 0.42 0.31 0.32 0.37 0.37 0.39
place 0.48 0.49 0.49 0.50 0.54 0.50 0.50 0.47 0.48 0.44 0.47 0.43 0.58 0.49 0.48 0.50 0.50 0.39 0.39 0.45 0.45 0.48
ADT decoration 0.28 0.28 0.26 0.26 0.27 0.26 0.26 0.26 0.24 0.27 0.23 0.31 0.25 0.28 0.28 0.28 0.26 0.15 0.12 0.20 0.14 0.25
meal 0.20 0.20 0.20 0.20 0.20 0.20 0.20 0.20 0.19 0.19 0.20 0.22 0.22 0.20 0.20 0.20 0.20 0.09 0.08 0.13 0.10 0.18
work 0.18 0.19 0.19 0.20 0.22 0.20 0.20 0.20 0.18 0.19 0.17 0.20 0.18 0.18 0.18 0.19 0.18 0.10 0.09 0.14 0.10 0.17
GIMO change 0.34 0.34 0.35 0.35 0.34 0.35 0.35 0.34 0.34 0.33 0.33 0.29 0.32 0.34 0.34 0.31 0.31 0.19 0.17 0.15 0.10 0.30
interact 0.38 0.38 0.37 0.37 0.36 0.37 0.37 0.36 0.36 0.35 0.36 0.32 0.36 0.38 0.38 0.35 0.34 0.21 0.21 0.18 0.15 0.33
rest 0.36 0.35 0.35 0.34 0.34 0.35 0.35 0.34 0.34 0.32 0.32 0.30 0.32 0.36 0.37 0.33 0.33 0.20 0.18 0.17 0.14 0.31
EgoBody catch 0.03 0.02 0.02 0.01 0.02 0.02 0.01 0.03 0.00 0.03 -0.02 0.04 0.00 0.03 0.02 0.02 0.02 0.02 0.00 0.02 0.01 0.02
chat 0.01 0.01 0.01 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.01 0.02 0.01 0.01 0.00 0.01 0.01 0.01 0.01 0.01 0.01
dance 0.05 0.05 0.05 0.04 0.04 0.04 0.05 0.04 0.04 0.03 0.04 0.03 0.03 0.05 0.05 0.05 0.04 0.02 0.01 0.02 0.02 0.04
discuss 0.02 0.02 0.03 0.03 0.04 0.03 0.03 0.04 0.03 0.02 0.03 0.03 0.03 0.01 0.01 0.00 0.02 0.00 0.00 0.01 0.01 0.02
learn 0.00 -0.01 -0.01 -0.01 -0.01 -0.01 0.00 -0.01 0.00 -0.01 -0.01 -0.01 0.00 0.00 0.00 -0.01 -0.01 0.00 0.01 0.01 0.01 0.00
perform 0.04 0.04 0.02 0.02 0.01 0.01 0.03 -0.01 0.01 -0.03 0.01 0.01 0.02 0.04 0.05 0.03 0.01 0.01 0.03 0.02 0.03 0.02
teach 0.00 0.00 0.01 0.01 0.02 0.01 0.01 0.00 0.02 0.00 0.02 0.01 0.02 0.00 0.01 0.01 0.02 0.02 0.02 0.02 0.02 0.01
To investigate any potential time delays between body motions and eye movements, we added different time intervals between body motions and gaze directions (time interval $>0$ means body motion data is moved forward in time while time interval $<0$ means gaze data is moved forward) and further calculated their correlations. Specifically, we first calculated the cosine similarities between eye gaze and the motions of different body joints and then averaged the cosine similarities across all the body joints. We can see from Figure 2 that in human-object interaction activities the highest correlations occur at between $400$ and $1500$ $ms$, suggesting that there exists a noticeable time delay between body motions and eye movements, i.e. eye movements precede body motions. In the human-human interaction activities, we find that the eye-body correlations are consistently small ($<$ 0.1 in all the cases) at different time intervals, validating that there is little or no correlation between eye gaze and body motion in human-human interactions.
(a) MoGaze
(b) ADT
(c) GIMO
(d) EgoBody